# Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

### v3

@@ Methods
Graph completion
In the analysis depicted in Figure 3, we enhance the traditional Jaccard similarity, hereafter referred to as $J$, to form a comprehensive similarity metric for evaluating potential material-to-application links. The Jaccard similarity is mathematically defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
where $A$ and $B$ represent the attribute sets of two compared materials, such as their properties or applications. To refine our predication method, we extend this basic formulation by incorporating weighted factors that reflect the relevance and impact of each attribute, resulting in a modified similarity score for "Materials-Applications", $J_{\text{mat}}$. This score is computed as follows:
@@
where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application. This approach mimics the investigative processes commonly employed by materials scientists when developing new materials. In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field. This enhanced method not only accounts for the quantity of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application. A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application. Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
Figure 3: The process of network-based graph completion.
In addition to the Enhanced Jaccard Similarity, we employ the TransE model [31]. TransE treats entities and relations as vectors in the same embedding space. The core idea of TransE is to model relations by interpreting them as translations in the embedding space. For a given relation $r$ between head $h$ and tail $t$, the relationship is said to hold if:
$\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|\approx 0$
[…]
Result
@@
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG series [2], [18], highlighting the differences in their construction methodologies as illustrated in Figure 4. However, the relations in MatKG is based on statistical means, which can not guarantee high accuracy of the relations, especially when the article contains multiple materials. Therefore, our main comparison object is MatKG2. MatKG2 is constructed using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in the Table 1. Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science. However, there is no marked difference in the performance of these models on ER tasks, we can say LLM has a weaker ability to complete ER in our task. This may be attributed to the limited contextual memory capabilities of the LLMs and this is we need to apply additional ER process. The normailzed Darwin shows the performance of our ER and normalization process, the result indicate that it not only achieve the ER task successfully, but also contribute to the NER and RE task.
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
@@
To better understand the enhancements provided by each component in the ER/Normalization process, ablation studies were performed. The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER. The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels which are included in the dictionary. Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExactor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED. Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications. ER-N/F is also an indispensable part of the ER process. The normalized data is transformed into triples to construct the MKG, which comprises 162,605 nodes and 731,772 edges, as illustrated in the schematic diagram in Figure 5.
@@
Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on publication year of paper. KG with earlier years applied similarity calculation algorithms to predict potential links, while KG with later years are used to verify how many sets of "Material Applications" in these predictions were reported in the following n years. The result shows in Figure 6. Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predication and the percentage of reported predication in next few year is displayed. The predictive capacity of the MKG is substantial, therefore, for each prediction, we only use the top 200 "material-application" pairs for validation. Looking at the results, as time progresses, the predicted materials are gradually verified. Using data from 2014, within nine years, 48.5% of the "material-application" predictions have been validated. This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields. The red, blue and green line are average percentage of reported link based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link predication, respectively. Network-based similarity has the highest accurate predication which indicates the effective of simulating the thinking of real material development using algorithms. The Figure 6 (b) is an example, visualizing the KG made from data before 2018, and the predicted "Material Application" reported within 5 years using network based algorithms.
@@
Figure 6: Validation of the graph completion. (a) Percentage of reported predication after years. (b) Example of predicated material - application using data before 2018.
Conclusion and future work
@@
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts. This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets. Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text. This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted. Moreover, this method can automatically extract high-precision knowledge form corpus, leading to dynamic KGs.
@@
Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis confirms the effectiveness and credibility of the MKG. We have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding, which enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions. First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG. This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature. Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG. Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science. Lastly, the integration of our Material Knowledge Graph with existing general material KGs, such as MatKG series and BoschKS, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
Appendix
A. Validation for MKG
@@
To validation of MKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 3. The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous normalization of entities under these categories, using our expert dictionary. However, strict normalization can also lead to certain entity loss, which is difficult to avoid. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory. "Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. Fortunately, the impact of this kind of error is not significant and will not bring a significant influence, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
@@
B. Predication result
As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023. It is important to note that these materials-application pairs were not ranked within the top five before 2017, which is why they do not appear in the figure for those earlier years. Stars mark the years these materials were featured in literature for their respective applications. The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas. This predication serves as an inspiration to materials researchers, accelerating progress in material development. For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER). This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
C. Paramaters for LLM fine tuning
epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0. warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: "LlamaDecoderLayer" token max length: 1024
Figure 7: High rank predicated materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, the stars represent the material has been reported in this application.
D. Experiments compute resources
The compute resources were primarily consumed in LLM fine-tuning, involving the use of 8 GPUs for 9.02 hours with a 36% utilization rate, consuming 25.98 service units and using 14245 MiB of storage from a quota of 81920 MiB.

### v5

@@ 2.4 Knowledge graph construction
2.5 Graph completion
The process of Graph Completion (GC) is shown in Figure 3 (a), where we perform link prediction through GC, segmented into four primary stages: graph splitting, similarity calculation, validation and evaluation, and parameter optimization. This structured approach ensures a comprehensive exploration of the link prediction capabilities. Graph splitting is meticulously performed based on the chronological assignment of the nodes. Nodes encapsulated within a defined prediction window $\delta$ are utilized as $G_{\text{ver}}$, which are designated for the validation of predictions. Nodes preceding this predictive interval serve as $G_{\text{tra}}$, employed to train and refine the prediction models. During the similarity calculation stage, advanced graph algorithms and embeddings are applied on $G_{\text{tra}}$ to ascertain the similarity between the "Materials" and "Applications" nodes. This phase leverages an enhanced Jaccard similarity metric, partitioned into three distinct components: $S(m,a)$, $F(m,a)$, and $T(m,a)$, which collectively improve the specificity of predictions.
Figure 3: (a)The process of MKG graph completion and (b) the schematic diagram of nodes comparison.
In the validation and evaluation phase, the sorted "Material" and "Application" nodes undergo rigorous testing on $G_{\text{ver}}$. The effectiveness of the predictive model is assessed by tallying the incorrect predictions, represented as En, which serves as a crucial measure of performance. The final stage, parameter optimization, involves the meticulous adjustment of the parameters $S(m,a)$, $F(m,a)$, and $T(m,a)$ based on their performance on $G_{\text{ver}}$. This iterative process is crucial for minimizing the count of En and enhancing the overall accuracy of the GC method. This adaptive adjustment underscores the dynamic and responsive nature of the model in refining link prediction accuracy within the graph completion framework.
It is worth mentioning that from a global perspective, we not only need to consider whether an "Application" may be similar to a "Material" node. We also need to consider whether the "Material" currently associated with this "Application" is similar to the potential "Material". Therefore, further improvement of Jaccard Similarity:
@@ 2.5 Graph completion
where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application. As shown in Figure 3 (b), this approach mimics the investigative processes commonly employed by materials scientists when developing new materials. In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field. This enhanced method not only accounts for the number of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application. A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application. Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
In addition to the Enhanced Jaccard Similarity, we employ the TransE model [bordes2013translating]. TransE treats entities and relations as vectors in the same embedding space. The core idea of TransE is to model relations by interpreting them as translations in the embedding space.
3 Result
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [MatKG2], highlighting the differences in their construction methodologies as illustrated in Figure 4. MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
@@ 3 Result
To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in Table 1. Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science. However, there is no marked difference in the performance of these models on ER tasks; we can say LLM has a weaker ability to complete ER in our task. This may be attributed to the limited contextual memory capabilities of the LLMs, and this is why we need to apply additional ER processes. The normailzed Darwin shows the performance of our ER and normalization process, the result indicates that it not only achieves the ER task successfully, but also contributes to the NER and RE task.
Table 1: Result of NER, RE, and ER through Fine-tuned LLMs.
@@
To better understand the enhancements provided by each component in the ER/Normalization process, ablation studies were performed. The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER. The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels that are included in the dictionary. Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExtractor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED. Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications. ER-N/F is also an indispensable part of the ER process. The normalized data is transformed into triples to construct the MKG, which comprises 162,605 nodes and 731,772 edges, as illustrated in the schematic diagram in Figure 5.
@@
Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on the paper’s publication year. KG, with earlier years, applied similarity calculation algorithms to predict potential links, while KG, with later years, is used to verify how many sets of "Material-Applications" in these predictions were reported in the following n years. The result is shown in Figure 6. Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predictions, and the percentage of reported predictions in the next few years is displayed. The predictive capacity of the MKG is substantial; therefore, for each prediction, we only use the top 200 "material-application" pairs for validation. Looking at the results, as time progresses, the predicted materials are gradually verified. Using data from 2014, within nine years, 48.5% of the "material-application" predictions have been validated. This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields. The red, blue, and green lines are the average percentage of reported links based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link prediction, respectively. Network-based similarity has the highest accurate prediction, which indicates the effectiveness of simulating the thinking of real material development using algorithms. The Figure 6 (b) is an example, visualizing the KG made from data before 2018 and the predicted "Material Application" reported within 5 years using network-based algorithms.
@@
Figure 6: Validation of the graph completion. (a) Percentage of reported prediction after years. (b) Example of predicted material - application using data before 2018.
4 Conclusion, Discussion, and Future Work
@@
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction, designed to efficiently extract triples from unstructured scientific texts. This methodology uniquely enables the fine-tuning of Large Language Models (LLMs) using minimally annotated datasets. The fine-tuned LLMs are then utilized to extract structured information from extensive corpora of unstructured text, bypassing predictive modeling to enhance the authenticity and traceability of the structured data extracted.
@@
Utilizing this approach, we constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities, such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis not only confirms the effectiveness and credibility of the MKG but also enriches our understanding of the material science domain.
Additionally, our methodology and the resulting KG hold significant potential across various dimensions:
  1. 1.
    Currently, our original text consists of article abstracts, which may omit a substantial amount of experimental methods and data. Fortunately, our method can be directly applied to full-text extraction, allowing us to more accurately capture subtle distinctions within complex scientific literature.
  2. 2.
    The ontology design of MKG has improved the precision of entity labeling in our pipeline and facilitates more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG.
  3. 3.
    The adaptability of our NLP pipeline to a broad range of scientific fields suggests its potential as a foundational template for constructing domain-specific KGs beyond materials science.
  4. 4.
    Integrating our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
For future work, given the apparent predictive capacity of our method for link prediction, we plan to focus on:
  1. 1.
    Analyzing Historical Patterns: By incorporating additional attributes such as authorship, publication year, and affiliations into our Knowledge Graph, we focus on analyzing historical "social" patterns. For example, investigating whether specific groups or institutions consistently engage in the repurposing of materials. This analysis will allow us to understand how the dynamics of material repurposing are influenced by collaborative networks or institutional trends. We will assess which materials or types of materials are more frequently repurposed and explore the potential social dynamics that drive these trends. This approach aims to provide insights into the lifecycle and innovation patterns of materials based on past activities and collaborations.
  2. 2.
    Identifying Future Materials and Applications: After constructing a complete MKG, in-depth graph neural networks and algorithms can be developed and used to explore potential relations in the graph. By exploring the data within our enriched MKG, we will identify new potential applications for the most promising materials or explore the future trend of materials. This proactive approach seeks to unlock new uses for materials before they become mainstream, providing a competitive edge in materials science. Additionally, by incorporating time dynamics, we aim to enhance the utility of our pipeline, allowing for time-aware graph embedding techniques for graph completion. This will enable us to explore how material properties and applications evolve over time, providing a dynamic perspective on the development and use of materials in various fields.
  3. 3.
    Studying Cluster Formation: We plan to analyze the formation of clusters within the KG, which are based on the links between materials. Understanding these clusters will help reveal how materials are interconnected, which can aid in discovering the connections between different materials that were not apparent from isolated studies.
5 Acknowledgments
This study was undertaken with the assistance of resources and services from the GreenDynamics and National Computational Infrastructure (NCl), which is supported by the Australian Government, the department of education, and through the UNSW-NCl partner trial scheme. Microsoft Research also provides resources and support to this study through the Acceleration Foundational Models Research Grant. We thank UNSW Library for assisting our team in obtaining text and data mining agreements with publishers.
[…]
Appendix A Appendix
A.1 Validation for MKG
@@
To validate MKG, we extract the 500 triples that are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, the annotator can obtain 1000 entities and 500 relations. The report from the annotator is shown in Table 3. The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous normalization of entities under these categories using our expert dictionary. However, strict normalization can also lead to certain entity loss, which is difficult to avoid. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, apply a less stringent standardization process, which leads to a certain degree of imprecision in entity and relation. But the results are still satisfactory. "Name" and "Acronym" have also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. Fortunately, the impact of this kind of error is not significant and will not bring a significant influence, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
@@
A.2 prediction result
As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023. It is important to note that these materials-application pairs were not ranked within the top five before 2017, which is why they do not appear in the figure for those earlier years. Stars mark the years these materials were featured in literature for their respective applications. The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas. This prediction serves as an inspiration to materials researchers, accelerating progress in material development. For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER). This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
A.3 Paramaters for LLM fine tuning
epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0. warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: ’LlamaDecoderLayer’ token max length: 1024
Figure 7: Top five potential materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, predicted on all available data (2014 - 2023) and their rank trends in the past years (2014 - related year). The stars represent the material has been reported in the related year.
A.4 Experiments compute resources
The compute resources were primarily consumed in LLM fine-tuning, which involved using 8 GPUs for 9.02 hours with a 36% utilization rate, 25.98 service units, and 14245 MiB of storage from a quota of 81920 MiB.